OpenStudy (anonymous):

Find the exact value of sin ( 17 pi / 12)

OpenStudy (anonymous):

Just need a little help

OpenStudy (cggurumanjunath):

http://www.youtube.com/watch?v=u12s3W0mk04

OpenStudy (anonymous):

hint: 17 pi / 12 = 12 pi /12 + 3 pi / 12 + 2 pi / 12 sin (3 pi / 12) = sin (pi /4) sin (2 pi / 12) = sin (pi / 6)

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

Amazing. There's a youtube video for just about anything you can think of now!